a chem lab keeps 2 acids solutions on hand. one is 20% acid and the other is 35% acid. How much 20% and how much 35% acid solution should be used to prepare 25 liters of a 26% acid solution?
Answer by stanbon(75887)

a chem lab keeps 2 acids solutions on hand. one is 20% acid and the other is 35% acid. How much 20% and how much 35% acid solution should be used to prepare 25 liters of a 26% acid solution?
Equations:
Quantity Eq.: x + y = 25
Acid Eq::::0.20x + 0.35y = 0.26*25
------
Multiply thru the Quantity Eq. by 20
Multiply thru the Acid Eq by 100
----------------
20x + 20y = 20*25
20x + 35y = 26*25
----
Subtract 1st from 2nd to get:
15y = 6*25
y = 10 liters (amt. of 35% solution in the mixture.
-----
x = 25-y = 15 liters (amt. 20% solution in the mixture.
========================================================
Cheers,
Stan H.